# Symmetric Iterations with Countable Support: A Framework for Choiceless ZF Extensions

### v1

@@ Abstract
We develop a framework for iterated symmetric extensions with countable support. At limit stages of uncountable cofinality we use the direct-limit presentation; at limits of cofinality $\omega$ we use a countable-support (inverse-limit) presentation via trees of conditions along a fixed cofinal sequence. We define the associated limit symmetry filter by head pushforwards and then (only at $\mathrm{cf}(\lambda)=\omega$) closing under countable intersections in the minimal normal way, and we prove that the resulting limit filter is normal and $\omega_{1}$–complete. This yields closure of hereditarily symmetric names under countable tuples and, consequently, preservation of $\mathrm{ZF}$; in a $\mathrm{ZFC}$ ground model we also obtain $\mathrm{DC}$ in the resulting symmetric model. When the iteration template is first-order definable over a $\mathrm{GBC}$ ground with Global Choice and sufficient class-recursion, the same scheme extends to class-length iterations, with the final symmetric model obtained from class hereditarily symmetric names (rather than as a union of stage models).
@@ 1 Introduction
Symmetric extensions are a robust method for producing models of $\mathrm{ZF}$ in which the Axiom of Choice fails while retaining strong combinatorial structure. While finite-support symmetric iterations are well understood (Karagila [1]), moving to countable support requires different machinery: filters must be genuinely complete (beyond finite intersections), and limit stages must be organized so that symmetry is preserved across many coordinates.
Aim.
We develop a uniform framework for iterated symmetric extensions with countable support. Set-length iterations are handled uniformly, and—when the iterand template is first-order definable over a Gödel–Bernays set theory with Global Choice ground—class-length iterations are admitted as well (Theorem 3.27).
Structure.
Section 2 fixes preliminaries and notation. Section 3 develops countable-support iterations (limit filters, HS names, $\mathrm{ZF}/\mathrm{DC}$ preservation). A brief discussion of examples is deferred to a companion application paper.
Scope of results.
For countable support, the direct/inverse limit presentations yield an $\omega_{1}$-complete limit filter and thus closure of HS names, $\mathrm{ZF}$, and $\mathrm{DC}$. Concrete applications require specifying successor-stage symmetric systems and verifying the intended choiceless consequences.
Why countable support?
Finite-support methods only require finite-intersection closure in the symmetry filter, which is too weak at limits of cofinality $\omega$. Countable support forces $\omega_{1}$-completeness of the limit filter, exactly the strength needed for Replacement and for preserving $\mathrm{DC}$. We isolate the finite-support failure points abstractly in Remark 4.1.
Preservation mechanism vs. productive conditions.
Our route to $\mathrm{ZF}+\mathrm{DC}$ uses filter completeness (plus stage-bounding and factorization); it complements finite-support, productive iteration techniques by supplying the closure needed in the infinite-support regime where finite-block tracking does not directly apply.
Terminology.
“Symmetric iteration” here means an iteration of symmetric systems with automorphism groups and normal filters of subgroups; “HS names” are hereditarily symmetric names with respect to the induced limit filter.
2 Preliminaries and Definitions
@@
Unless explicitly stated otherwise, we carry out the iteration construction and the bookkeeping about supports, tuples of names, and filter generation in a fixed ambient background universe satisfying $\mathrm{ZFC}$ (and, for the class-length development in §3.6, in a ground model of $\mathrm{GBC}$ with Global Choice). The purpose of the paper is not to assume Choice in the symmetric models produced: our preservation theorems establish that the resulting symmetric submodels satisfy $\mathrm{ZF}$ (and specified fragments of Dependent Choice), while $\mathrm{AC}$ typically fails there (depending on the iterated forcing notion).
When we say a support is countable, we mean countable in the ambient background universe in which the iteration is carried out. We do not assert that these supports are countable in intermediate symmetric submodels.
@@
A symmetric system is a triple $(\mathbb{P},\mathcal{G},\mathcal{F})$ where:
  • $\mathbb{P}$ is a notion of forcing;
  • $\mathcal{G}$ is a group of automorphisms of $\mathbb{P}$;
  • $\mathcal{F}$ is a normal filter of subgroups of $\mathcal{G}$.
Remark 2.3 (Tenacity and excellent supports; compatibility with [1]).
Karagila (with Hayut) shows that every symmetric system is equivalent, on a dense set, to a tenacious one (Appendix A, Thm. A.2 in [1]). We do not rely on tenacity per se—our arguments use normal filters and HS stabilizers (and, later, tuple–stabilizer compression)— but one may refine our presentations on a dense set to tenacious systems without affecting the proofs. Moreover, our “finite-generator” behavior (closure under finite intersections of basic generators) mirrors Karagila’s excellent supports and their stability under intersections (Def. 4.3 and Prop. 4.4 in [1]).
Convention 2.4 (Tenacity assumption).
By Appendix A of [1], every symmetric system is equivalent to a tenacious one, so we may assume tenacity without loss of generality when convenient. This aligns with DC-preservation criteria in the literature (cf. [4]).
Definition 2.5 (Hereditarily Symmetric Name).
A $\mathbb{P}$-name $\dot{x}$ is hereditarily symmetric (HS) if:
  • $\mathrm{sym}(\dot{x})\in\mathcal{F}$, where $\mathrm{sym}(\dot{x})$ is the group of automorphisms fixing $\dot{x}$;
  • All $\dot{y}\in\mathrm{dom}(\dot{x})$ are also HS.
Definition 2.6 (Dependent Choice $DC_{\mu}$).
For a nonzero ordinal $\mu$, $DC_{\mu}$ asserts: for every binary relation $R$ on a class $A$ with $\forall x\in A\,\exists y\in A\,(xRy)$, there exists a sequence $\langle a_{\xi}:\xi<\mu\rangle$ in $A$ such that $a_{\xi}Ra_{\xi+1}$ for all $\xi+1<\mu$.
Remark 2.7 (Limit-stage bookkeeping).
At limits we fix the following presentation once and for all: if $\mathrm{cf}(\lambda)>\omega$, use the direct limit $\mathbb{P}_{\lambda}=\bigcup_{\beta<\lambda}\mathbb{P}_{\beta}$. Since every countable $A\subseteq\lambda$ is bounded in $\lambda$ when $\mathrm{cf}(\lambda)>\omega$, this coincides with the usual countable-support limit at such $\lambda$ (and head pushforwards are just inclusions). The identification of initial segments and their tails uses the canonical projection and two-step factorization (Fact 2.8). If $\mathrm{cf}(\lambda)=\omega$, use the inverse limit along a fixed cofinal sequence. All later proofs refer to this convention.
Notation.
Throughout, supports are countable. We write $\mathrm{cf}(\lambda)$ for cofinality, $\mathrm{supp}(\dot{x})$ for the set of coordinates a name depends on, and $\mathrm{fix}_{\mathcal{G}}(\dot{x})=\{g\in\mathcal{G}:g(\dot{x})=\dot{x}\}$ for the stabilizer. For $\beta<\lambda$, $\pi^{\beta}_{\lambda}:\mathcal{G}_{\lambda}\to\mathcal{G}_{\beta}$ and $\pi^{\beta}_{\lambda}:\mathbb{P}_{\lambda}\to\mathbb{P}_{\beta}$ denote the canonical projections.
$\lambda$ iteration length (limit stages treated via Remark 2.7)
$\omega_{1}$ completeness level for limit filters (countable intersections)
$\mathbb{P}_{\lambda}$ iteration forcing up to stage $\lambda$
$\mathcal{G}_{\lambda}$ automorphism group at stage $\lambda$
$\mathcal{F}_{\beta}$, $\tilde{\mathcal{F}}_{\lambda}$ stage filter / limit filter
$\pi^{\beta}_{\lambda}$ projection on forcing/groups (context determines target)
$\mathrm{supp}(\dot{x})$ coordinates used by the name $\dot{x}$
$\mathrm{fix}(\dot{x})$ stabilizer of $\dot{x}$ in the relevant group
HS “hereditarily symmetric” (Definition 3.16)
Fact 2.8 (Two-step projection/factorization).
@@ A $\mathbb{P}$-name $\dot{x}$ is hereditarily symmetric (HS) if:
3 Countable-Support Symmetric Iterations
@@ 3.1 Framework and Setup
Definition 3.1 (Countable Support Symmetric Iteration).
A countable support symmetric iteration of length $\Theta$ is constructed recursively as follows:
Stage 0 (Base):
  • $\mathbb{P}_{0}=\{1_{0}\}$ (trivial forcing with one element)
  • $\mathcal{G}_{0}=\{\mathrm{id}\}$ (trivial automorphism group)
  • $\mathcal{F}_{0}=\{\mathcal{G}_{0}\}$ (trivial filter containing only the whole group)
Stage $\alpha=\beta+1$ (Successor):
Given $(\mathbb{P}_{\beta},\mathcal{G}_{\beta},\mathcal{F}_{\beta})$ and a $\mathbb{P}_{\beta}$-name $\dot{\mathbb{Q}}_{\beta}$ for a forcing notion:
  • Forcing: $\mathbb{P}_{\alpha}=\mathbb{P}_{\beta}*\dot{\mathbb{Q}}_{\beta}$ consists of pairs $(p,\dot{q})$ as usual.
  • Automorphisms: $\mathcal{G}_{\alpha}$ is generated by lifts from $\mathcal{G}_{\beta}$ and automorphisms of $\dot{\mathbb{Q}}_{\beta}$.
  • Filter: $\mathcal{F}_{\alpha}$ is a normal filter of subgroups of $\mathcal{G}_{\alpha}$ extending the pushed-forward copy of $\mathcal{F}_{\beta}$:
    1. 1.
      Pushforwards: for $H\in\mathcal{F}_{\beta}$ include the head pushforward
      $\pi^{\beta}_{\alpha,*}(H)\ :=\ \{g\in\mathcal{G}_{\alpha}:\pi^{\beta}_{\alpha}(g)\in H\},$
      where $\pi^{\beta}_{\alpha}:\mathcal{G}_{\alpha}\to\mathcal{G}_{\beta}$ is the canonical restriction (head projection).
    2. 2.
      Iterand symmetry generators: additional subgroups coming from the chosen symmetric system on $\dot{\mathbb{Q}}_{\beta}$.
    For the $\mathrm{DC}$-preservation theorem below, we will assume (and in applications verify) that each $\mathcal{F}_{\alpha}$ is $\omega_{1}$–complete.
Stage $\alpha=\lambda$ (Limit, $\mathrm{cf}(\lambda)=\omega$):
  • Forcing: $\mathbb{P}_{\lambda}$ is the usual countable-support limit of the iteration. A condition $p\in\mathbb{P}_{\lambda}$ is a partial function with countable domain $\operatorname{dom}(p)\subseteq\lambda$ such that for each $\beta\in\operatorname{dom}(p)$, $p(\beta)$ is a $\mathbb{P}_{\beta}$–name for a condition in $\dot{\mathbb{Q}}_{\beta}$. The order is the standard one: $q\leq p$ iff $\operatorname{dom}(q)\supseteq\operatorname{dom}(p)$ and for all $\beta\in\operatorname{dom}(p)$,
@@
    We define the support of $p$ by $\mathrm{supp}(p):=\operatorname{dom}(p)$, so $\mathrm{supp}(p)$ is countable by definition.

### v3

# Limit Filters and Dependent Choice in Countable-Support Symmetric Iterations

@@ Abstract
We isolate the limit-stage filter construction needed for countable-support symmetric iterations built from standard successor-step symmetric systems. At successor stages we take the $\omega_{1}$-completion of the usual successor-stage symmetry filter. At limit stages of uncountable cofinality, countable supports are bounded and the direct-limit filter is $\omega_{1}$-complete by stage-bounding; at limits of cofinality $\omega$, we define the limit filter as the smallest normal $\omega_{1}$-complete filter extending the head-pullback generators. In all cases the resulting limit filter is normal and $\omega_{1}$-complete.
Using these limit filters, we prove that the class of hereditarily symmetric names is closed under the operations required for $\mathrm{ZF}$, that the resulting symmetric model satisfies $\mathrm{ZF}$ (Theorem 3.17), and that over a $\mathrm{ZFC}$ ground it satisfies $\mathrm{DC}=\mathrm{DC}_{\omega}$ (Theorem 3.20). We also explain why no general class-length $\mathrm{ZF}$-preservation theorem is claimed here, although specific class-length symmetric iterations may still be analyzable separately over a $\mathrm{GBC}$ background.
As a self-contained application, we construct, for any uncountable cardinal $\kappa$ with $\mathrm{cf}(\kappa)\geq\omega_{1}$, a model of $\mathrm{ZF}+\mathrm{DC}+\neg\mathrm{AC}_{\kappa}(\mathcal{F})$, where $\mathcal{F}$ is a $\kappa$-indexed family of $2$-element sets of reals with no choice function. Each step of the iteration adds one unpickable pair, so the degree of $\mathrm{AC}$-failure is controlled by the iteration length. We also prove that the analogous finite-support construction fails $\mathrm{DC}$ at the first $\omega$-limit stage, showing why $\omega_{1}$-complete limit filters are structurally needed for this application.
@@ 1 Introduction
Symmetric extensions produce intermediate models of $\mathrm{ZF}$ between a ground model $V$ and a forcing extension $V[G]$ by restricting attention to the class of hereditarily symmetric names. Karagila developed a general framework for finite-support symmetric iterations [3]. Countable support introduces a new constraint at limit stages of cofinality $\omega$: to keep the class of hereditarily symmetric names closed under countable constructions, one needs limit-stage symmetry filters which are not merely normal but also $\omega_{1}$–complete.
What is proved here.
Working in a fixed background universe $V$, we assume the standard successor-step construction of symmetric systems (as in [3, §§3–5]) and isolate the additional limit-stage technology needed for countable-support iterations. At each limit $\lambda$ we define a canonical limit filter $\tilde{\mathcal{F}}_{\lambda}$ extending the head pullbacks of earlier-stage filters and prove:
  • $\tilde{\mathcal{F}}_{\lambda}$ is normal and $\omega_{1}$–complete (Theorem 3.13);
  • the class $\mathrm{HS}_{\lambda}$ of hereditarily symmetric names is closed under the operations required in the standard symmetric-model verification of $\mathrm{ZF}$ (Theorems 3.15 and 3.17);
  • if $V\models\mathrm{ZFC}$, then the resulting symmetric model satisfies $\mathrm{DC}=\mathrm{DC}_{\omega}$ (Theorem 3.20).
Why the limit filters matter.
The central point is that countable-support iterations require $\omega_{1}$-complete symmetry filters at limit stages. This has two consequences needed in the present paper. First, $\omega_{1}$-completeness ensures that hereditarily symmetric names are closed under countable tuples (Theorem 3.15), yielding $\mathrm{DC}$ in the resulting model (Theorem 3.20) and ruling out Dedekind-finite or amorphous pathologies (Corollary 3.22). Second, the same limit-filter structure is what makes the later application possible: the iterated unordered-pairs construction of Section 4 uses these filters to build, for each uncountable $\kappa$ with $\mathrm{cf}(\kappa)\geq\omega_{1}$, a model of $\mathrm{ZF}+\mathrm{DC}+\neg\mathrm{AC}_{\kappa}(\mathcal{F})$.
A finite-support version of the same step forcing fails to produce a $\mathrm{DC}$ model at the $\omega$-th limit stage (Proposition 4.4), showing that the $\omega_{1}$-complete limit-filter construction is not merely convenient but structurally necessary for this application.
What is not claimed.
We do not claim here a complete general theory of countable-support symmetric iterations comparable in scope to broader support frameworks. The successor-step construction is taken from the standard theory, and the new contribution of the present paper is the limit-stage filter technology needed for countable support together with the resulting $\mathrm{ZF}/\mathrm{DC}$ preservation theorems.
We also do not claim that the framework by itself produces new choiceless consequences; those arise from the specific iteration templates used in applications (cf. Section 4). In particular, we do not claim any uncountable-length dependent choice (e.g. $\mathrm{DC}_{\omega_{1}}$) without additional hypotheses, and we do not address cardinal preservation or other fine structure beyond what is explicitly proved. All uses of “countable” refer to countability in the background universe $V$ (Remark 2.1); for the only countability closure needed here (unions of countable supports, i.e. subsets of ordinals), see Remark 2.2.
Organization.
Section 2 fixes forcing/symmetry notation and records the Symmetry Lemma and related invariance facts used later. Section 3 develops the countable-support iteration, the limit filters, and the closure and preservation theorems. Remark 3.25 isolates the points where finite support is insufficient for the present arguments. Section 4 constructs the iterated unordered pairs model as a self-contained application, producing $\mathrm{ZF}+\mathrm{DC}+\neg\mathrm{AC}_{\kappa}(\mathcal{F})$ for any uncountable $\kappa$ with $\mathrm{cf}(\kappa)\geq\omega_{1}$, and proves that the finite-support version of the same step forcing fails to preserve $\mathrm{DC}$.
2 Preliminaries: forcing, automorphisms, and symmetric names
2.1 Background theory and metatheoretic conventions
@@
All constructions of forcing notions, names, automorphism groups, filters, and supports are carried out in a fixed background universe $V$.
  • For the set-length development (existence of the iteration, the limit filters, and $\mathrm{ZF}$ preservation) we work in $V\models\mathsf{ZF}$.
  • For the dependent-choice conclusion (Theorem 3.20) we additionally assume $V\models\mathsf{ZFC}$.
When we say a set is countable, this is always in the sense of the background universe $V$ (equivalently: there is an injection into $\omega$ in $V$).
No claim is made that these sets are countable inside intermediate symmetric models.
Remark 2.2 (Countable unions of countable supports in $\mathrm{ZF}$).
All countability bookkeeping in Sections 33.5 concerns subsets of ordinals (supports and stage indices). In $\mathrm{ZF}$, if $\langle A_{n}:n<\omega\rangle$ is a sequence of countable subsets of an ordinal $\lambda$, then $\bigcup_{n<\omega}A_{n}$ is countable. Indeed, each $A_{n}\subseteq\lambda$ admits a canonical increasing enumeration (defined by recursion using the well-ordering of $\lambda$, padded by 0 if $A_{n}$ is finite), and then $(n,m)\mapsto e_{n}(m)$ is a surjection from $\omega\times\omega$ onto $\bigcup_{n<\omega}A_{n}$. Thus no separate choice principle about countable unions of arbitrary sets is used for the set-length iteration technology here.
2.2 Forcing and names
We use standard forcing conventions. A forcing notion is a poset $\mathbb{P}=(P,\leq)$ with top condition $1_{\mathbb{P}}$. A $\mathbb{P}$-name is a set $\dot{x}$ of pairs $\langle\dot{y},p\rangle$ with $\dot{y}$ a $\mathbb{P}$-name and $p\in\mathbb{P}$. For a $V$-generic filter $G\subseteq\mathbb{P}$, $\dot{x}^{G}$ denotes the usual valuation.
2.3 Automorphisms and symmetric systems
Definition 2.3 (Automorphisms).
An automorphism of $\mathbb{P}$ is a bijection $\pi:P\to P$ satisfying $p\leq q\iff\pi(p)\leq\pi(q)$ and $\pi(1_{\mathbb{P}})=1_{\mathbb{P}}$. If $\mathcal{G}\leq\mathrm{Aut}(\mathbb{P})$ is a subgroup, we write $\pi\cdot p$ for $\pi(p)$.
Definition 2.4 (Normal filter of subgroups).
Let $\mathcal{G}$ be a group. A filter of subgroups $\mathcal{F}$ on $\mathcal{G}$ is a nonempty collection of subgroups of $\mathcal{G}$ such that:
  1. 1.
    $\mathcal{G}\in\mathcal{F}$;
  2. 2.
    if $H\in\mathcal{F}$ and $H\leq K\leq\mathcal{G}$, then $K\in\mathcal{F}$;
  3. 3.
    if $H,K\in\mathcal{F}$, then $H\cap K\in\mathcal{F}$.
It is normal if for all $H\in\mathcal{F}$ and all $\pi\in\mathcal{G}$, $\pi H\pi^{-1}\in\mathcal{F}$. It is $\omega_{1}$-complete if $\bigcap_{n<\omega}H_{n}\in\mathcal{F}$ whenever $H_{n}\in\mathcal{F}$ for all $n<\omega$.
Definition 2.5 ($\omega_{1}$-completion of a normal filter).
Let $\mathcal{F}$ be a normal filter of subgroups on $\mathcal{G}$. Define its $\omega_{1}$-completion $\mathcal{F}^{\omega_{1}}$ by
$K\in\mathcal{F}^{\omega_{1}}\quad\Longleftrightarrow\quad\exists\langle H_{n}:n<\omega\rangle\in\mathcal{F}^{\omega}\ \Bigl(\ \bigcap_{n<\omega}H_{n}\leq K\ \Bigr).$
Lemma 2.6 ($\omega_{1}$-completion is the smallest normal $\omega_{1}$-complete extension).
If $\mathcal{F}$ is a normal filter of subgroups on $\mathcal{G}$, then $\mathcal{F}^{\omega_{1}}$ is a normal, $\omega_{1}$-complete filter of subgroups on $\mathcal{G}$, and it is the smallest normal, $\omega_{1}$-complete filter extending $\mathcal{F}$.
Proof.
Upward closure and containing $\mathcal{G}$ are immediate. $\omega_{1}$-completeness holds because if $K_{m}\in\mathcal{F}^{\omega_{1}}$ is witnessed by $\langle H^{m}_{n}:n<\omega\rangle\subseteq\mathcal{F}$, then $\bigcap_{m}K_{m}$ is witnessed by the single $\mathcal{F}$-sequence obtained by concatenating the sequences $\langle H^{m}_{n}:n<\omega\rangle$ over $m<\omega$. Normality is preserved since for $\pi\in\mathcal{G}$,
$\pi\Bigl(\bigcap_{n<\omega}H_{n}\Bigr)\pi^{-1}=\bigcap_{n<\omega}\pi H_{n}\pi^{-1},$
and $\mathcal{F}$ is normal. Minimality is immediate: any normal $\omega_{1}$-complete filter extending $\mathcal{F}$ must contain every countable intersection of members of $\mathcal{F}$, hence must contain $\mathcal{F}^{\omega_{1}}$. ∎
@@ 2.3 Automorphisms and symmetric systems
A symmetric system is a triple $\langle\mathbb{P},\mathcal{G},\mathcal{F}\rangle$ where $\mathbb{P}$ is a forcing notion, $\mathcal{G}\leq\mathrm{Aut}(\mathbb{P})$, and $\mathcal{F}$ is a normal filter of subgroups of $\mathcal{G}$.
2.4 Action on names and the Symmetry Lemma
Given $\langle\mathbb{P},\mathcal{G},\mathcal{F}\rangle$, define recursively the action of $\pi\in\mathcal{G}$ on $\mathbb{P}$-names by
$\pi\dot{x}\ :=\ \{\langle\pi\dot{y},\ \pi p\rangle:\langle\dot{y},p\rangle\in\dot{x}\}.$
Definition 2.8 (Stabilizers).
For a name $\dot{x}$, its (setwise) stabilizer is
$\operatorname{sym}_{\mathcal{G}}(\dot{x})\ :=\ \{\pi\in\mathcal{G}:\pi\dot{x}=\dot{x}\}.$
When $\mathcal{G}$ is clear from context, we write $\operatorname{sym}(\dot{x})$.
Lemma 2.9 (Symmetry Lemma / Permutation Lemma).
For any formula $\varphi$ in the forcing language, any $\pi\in\mathcal{G}$, any $p\in\mathbb{P}$, and any names $\dot{x}_{1},\dots,\dot{x}_{n}$,
$p\Vdash_{\mathbb{P}}\varphi(\dot{x}_{1},\dots,\dot{x}_{n})\quad\Longleftrightarrow\quad\pi p\Vdash_{\mathbb{P}}\varphi(\pi\dot{x}_{1},\dots,\pi\dot{x}_{n}).$
See, e.g., [3, Lemma 2.4].
2.5 Symmetric and hereditarily symmetric names
Definition 2.10 (Symmetric / hereditarily symmetric).
Let $\langle\mathbb{P},\mathcal{G},\mathcal{F}\rangle$ be a symmetric system. A name $\dot{x}$ is $\mathcal{F}$-symmetric if $\operatorname{sym}_{\mathcal{G}}(\dot{x})\in\mathcal{F}$. A name $\dot{x}$ is hereditarily $\mathcal{F}$-symmetric if $\dot{x}$ is $\mathcal{F}$-symmetric and every $\dot{y}$ appearing in the transitive closure of $\dot{x}$ (equivalently, every $\dot{y}\in\mathrm{dom}(\dot{x})$ and recursively) is hereditarily $\mathcal{F}$-symmetric. We write $\mathrm{HS}(\mathbb{P},\mathcal{G},\mathcal{F})$ for the class of hereditarily $\mathcal{F}$-symmetric $\mathbb{P}$-names.
2.6 Dependent Choice (set formulation)
Definition 2.11 (Dependent Choice).
Let $\mu$ be a nonzero ordinal. $\mathrm{DC}_{\mu}$ is the statement: for every set $A$ and every binary relation $R\subseteq A\times A$ such that $\forall x\in A\,\exists y\in A\ (xRy)$, there exists a sequence $\langle a_{\xi}:\xi<\mu\rangle$ of elements of $A$ with $a_{\xi}Ra_{\xi+1}$ for all $\xi+1<\mu$. We write $\mathrm{DC}=\mathrm{DC}_{\omega}$ and $\mathrm{DC}_{<\kappa}$ for $\forall\mu<\kappa\,\mathrm{DC}_{\mu}$.
2.7 Two-step iterations (factorization)
Fact 2.12 (Two-step factorization).
@@ 2.7 Two-step iterations (factorization)
Remark 2.13 (Notation discipline for later sections).
In the iteration sections we will not overload projection symbols. For forcing, we use restriction $p\upharpoonright\beta$ and (when needed) a map $\pi_{\beta,\lambda}:\mathbb{P}_{\lambda}\to\mathbb{P}_{\beta}$. For groups, we use a distinct restriction homomorphism $\rho_{\beta,\lambda}:\mathcal{G}_{\lambda}\to\mathcal{G}_{\beta}$.
3 Limit-Stage Filters for Countable-Support Symmetric Iterations
@@ 3.1 Framework and Setup
Definition 3.1 (Countable-support iteration setup).
Fix an ordinal $\Theta$. A countable-support symmetric iteration of length $\Theta$ is a system
together with, for each $\alpha<\Theta$, a $\mathbb{P}_{\alpha}$-name
$\dot{\mathscr{S}}_{\alpha}=\langle\dot{\mathbb{Q}}_{\alpha},\dot{\mathcal{H}}_{\alpha},\dot{\mathcal{K}}_{\alpha}\rangle$
such that
$1_{\mathbb{P}_{\alpha}}\Vdash_{\mathbb{P}_{\alpha}}\text{``$\dot{\mathscr{S}}_{\alpha}$ is a symmetric system.''}$
The recursion is as follows.
Stage 0.
$\mathbb{P}_{0}=\{1\}$, $\mathcal{G}_{0}=\{\mathrm{id}\}$, and $\mathcal{F}_{0}=\{\mathcal{G}_{0}\}$.
Successor stage $\alpha+1$.
Let $\dot{\mathscr{S}}_{\alpha}=\langle\dot{\mathbb{Q}}_{\alpha},\dot{\mathcal{H}}_{\alpha},\dot{\mathcal{K}}_{\alpha}\rangle$ be as above. Define
$\mathbb{P}_{\alpha+1}=\mathbb{P}_{\alpha}*\dot{\mathbb{Q}}_{\alpha}.$
Define $\mathcal{G}_{\alpha+1}$ and an intermediate normal successor-stage filter $\mathcal{F}^{\mathrm{Kar}}_{\alpha+1}$ by the standard two-step symmetric-system construction from $(\mathbb{P}_{\alpha},\mathcal{G}_{\alpha},\mathcal{F}_{\alpha})$ and the name $\dot{\mathscr{S}}_{\alpha}$: there is a canonical restriction homomorphism (“head restriction”)
$\rho_{\alpha,\alpha+1}:\mathcal{G}_{\alpha+1}\to\mathcal{G}_{\alpha}$
and $\mathcal{F}^{\mathrm{Kar}}_{\alpha+1}$ is a normal filter of subgroups of $\mathcal{G}_{\alpha+1}$ extending the pullbacks $\rho_{\alpha,\alpha+1}^{-1}(H)$ for $H\in\mathcal{F}_{\alpha}$ and the tail symmetries coming from $\dot{\mathcal{K}}_{\alpha}$. (We do not reprove this successor-step construction here; see, e.g., [3, §§3–5].)
For the countable-support limit-stage arguments below, we require $\omega_{1}$–completeness at intermediate stages, so we define
$\mathcal{F}_{\alpha+1}\ :=\ \bigl(\mathcal{F}^{\mathrm{Kar}}_{\alpha+1}\bigr)^{\omega_{1}},$
the $\omega_{1}$–completion of $\mathcal{F}^{\mathrm{Kar}}_{\alpha+1}$ in the sense of Definition 2.5 (which is the smallest normal $\omega_{1}$–complete extension by Lemma 2.6).
Limit stage $\lambda\leq\Theta$ with $\mathrm{cf}(\lambda)=\omega$.
$\mathbb{P}_{\lambda}$ is the usual countable-support limit of $\langle\mathbb{P}_{\alpha},\dot{\mathbb{Q}}_{\alpha}:\alpha<\lambda\rangle$: conditions are partial functions $p$ with countable domain $\operatorname{dom}(p)\subseteq\lambda$, where for each $\beta\in\operatorname{dom}(p)$, $p(\beta)$ is a $\mathbb{P}_{\beta}$-name for a condition in $\dot{\mathbb{Q}}_{\beta}$, ordered by $q\leq p$ iff $\operatorname{dom}(q)\supseteq\operatorname{dom}(p)$ and for all $\beta\in\operatorname{dom}(p)$,
@@
Define $\mathrm{supp}(p):=\operatorname{dom}(p)$.